FEL Upgrade Project Weekly Brief
June 26-30, 2000


This week the FEL was operated for two users and three different experiments involving pulse laser ablation, pulsed laser deposition and IR microscopy.


The project costbook was decreased by $276k from the baseline budget of $9.3M to account for a DoE facility tax added to the project. The scope decrease was taken in one area of cryomodule development which will be covered in the second phase of the project. All other accounts remain as presented at the project kick-off meeting on June 12-13. Since next week is a short work week with the July 4th holiday (even FEL folks take an occaisional day off...), there will no weekly brief
next Friday.

Design/Construction Activities:

Beam Physics: (WBS 3)

We received a proposal from the Naval Postgraduate School for modelling efforts that support the upgrade. The NPS provided valuable input to both the original IR Demo and the upgrade design and we welcome their new proposal.

Injector: (WBS 4)

We have finally been able to continue our qualification bake-out of the Apertured Cesiator after replacing both the ion pump and its controller. Testing should start next week after completion of the bake.We are planning an early test of the high voltage power supply for the gun to monitor its performance at 10 mA. Tentatively, we are scheduling this to follow the completion of the summer user run in August.

RF Systems: (WBS 6)

The first installation of waveguide components in the 3rd cryomodule position will occur early on the morning of July 10
prior to start of the first shift user run.

Instrumentation and Controls: (WBS 8)

Parts for prototype 3" BPM are nearly complete, the RF feedthrus are due Aug. 6, 2000. These will be "dropped" in as soon as they arrive, and then testing will begin. The drawings for the 3" shielded beam viewers are in detailing. They should be ready to send out for budgetary quotes in approximately two weeks. Progress is being made on new optical diagnostics; 2 pcs. 10mm square LiTa detectors with built-in preamps are in final testing. These coupled with sample and holds will be used as permanent FEL power meters and will display energy-per-micropulse.

Additionally quadrant LiTa are being assembled for use as an optical beam position monitoring system; this "box" will output volts/mm. All LSS systems have been recertified for the run. Status lamps have been added to each of the user labs to indicate state of the FEL - these are intended as idiot lights for the users, in the last run they had no simple indication of a failure.

Beam Transport: (WBS 10)

Dipoles - We continued layout of the GX Prototype, refining the details of the insulation system by consulting with vendors before freezing the configuration. This refinement changed the dimensions of the required steel slabs. We hope to order the prototype steel and copper next week. The engineer assigned to make the benchmark 3D magnetic model of an existing H-style dipole has been delayed due to pressing demands from existing projects. 3-inch quadrupole yoke machining has started and the annealing specification was signed off. Coil prints are being back checked.  Continued to create the 3-D magnetic model (TOSCA) of the 3-inch quadrupole. We have delayed making further samples of dipole chambers until the definition of the prototype dipoles is more advanced.

Wiggler (WBS 10)

Created the details of the Optical Klystron on its strong back with all diagnostics, vacuum pipe, dispersion section and water manifolds so that existing parts could be modified and new parts made. The bus bars and mounting plates were received.  The shop started making tooling in anticipation of making the coils. The insulated conductor was received from the insulation vendor. Finished the plug concepts for disassembly of electrical leads for power and safety circuits during installation.

Optics (WBS 11)

We have completed a preliminary layout of the optical cavity for combination with the electron beam transport lattice so that we identify interferences and interfaces. We have brought new optical diagnostics on line this week to support operations. There is now an on-line diagnostic for measurement of the FEL mode quality. We have received new 95% beamsplitters that can be added to the mirror cassets in the user labs to enhance flexibility for users.

Operations/ Commissioning

This week, the first full week of FEL operations for the summer user run, has seen more than typical fault rates with RF equipment, including a failed HV cable. Operations have been hampered also by investigations into the cause of an elliptical mode pattern in the optical mode in lab 6. This investigation required the removal and reinstallation of the cavity mirrors for inspection (they were fine), and explorations of different mirror and electron beam steering configurations to see where
the problem is arising. Analysis of the data to date indicates that one of the mirrors in the collimator can is warped. Despite these delays, data was taken on ablation rates, trepanning experiments were carried out, and the first scanning-near-field microscope images were acquired.

The laser has lased well, with over 1% efficiency obtained on a routine basis with 6% duty cycle at 3.1 microns. Mirrors have been received for operation at 1micron in July.

Two PbSe detectors have been built for the Norfolk State IR microscope experiment. These included servo control for the TE coolers internal to the detector.